passes through the point (1,-7) and is parallel to the line y=-2x+6
Answer by fractalier(6550)

If it is parallel to the given line, they have the same slope, -2.
You can plug everything into y = mx + b and find b...
y = mx + b
-7 = (-2)(1) + b
-7 = -2 + b
b = -5
and so we get
y = -2x - 5